pleasehelp!
solve for x: 8/x = 12/(x+2)
A. 4/5
B 1/2
C 4
Thank you!

Answer by checkley71(8403)

8/X=12/(X+2) OR 8X+16=12X OR 16=4X OR X=16/4 OR X=4